# The genealogy of nearly critical branching processes in varying environment

### v3

@@ 7 Convergence of the BPVE
This last section is dedicated to the proof of Theorem 2. Recall the encoding of generation $tN$ of the BPVE as a random metric measure space $[T_{tN},d_{tN},\lambda_{tN}]$ from Section 7.
As alluded to in the introduction, we follow the general proof strategy proposed in an earlier work by two of the authors [15]. We start by proving that the genealogy of the BPVE converges in the Gromov-weak topology. According to Proposition 3.1, proving convergence in distribution for the Gromov-weak topology amounts to computing the limit of the moments of the random metric measure space $[T_{tN},d_{tN},\lambda_{tN}]$. Using the many-to-few formula, see Proposition 5.3, these moments can be expressed in terms of the distribution of a finite tree, the $k$-spine tree [27, 26, 15]. This $k$-spine tree is studied in Section 7.1 where we derive the limit of the moments of the BPVE. A small caveat to this approach is that, for the moments to be well-defined, we need to make a truncation of the reproduction laws, which is carried out in Section 7.2. We relate the truncated and the untruncated versions via Corollary 3.4. Once the convergence is established in the Gromov-weak topology, we reinforce it to a convergence in the GHP topology by using Proposition 3.6.
@@ 7.1 Moments of the BPVE
Consider a BPVE satisfying (6) and (7), and a fixed $t>0$ such that $t$ is a continuity point of $\sigma^{2}$ and $X$, and $\sigma^{2}$ is strictly increasing in $t$. If the moment growth assumption (50) is fulfilled, then for any bounded continuous $\phi$,
@@
where $\sigma$ is an independent and uniform permutation of $[k]$ and with $\big{(}(H_{i,j})_{i,j}\big{)}$ as in (18).
@@
Let $|B_{u_{1},\dots,u_{k}}|$ denote the number of branch points of the subtree spanned by $(u_{1},\dots,u_{k})$. We decompose the expectation with respect to the value of $|B_{u_{1},\dots,u_{k}}|$
@@
There are two steps in the proof. First, we prove that, due to our assumption on the moments (50), the contributions of the terms with $b<k-1$ vanish. Secondly, we compute the limit of the term with only binary branch points and show that we recover the moments of the environmental CPP. Step 1. Fix $b<k-1$. We apply the many-to-few formula with branch lengths $H_{i}^{(N)}$ uniformly distributed on $\{1,\dots,tN\}$. Let $\mathbb{Q}^{k,Nt}$ be the corresponding $k$-spine tree distribution. By Proposition 5.3, we have
@@
where we have used Lemma A.1 in the second line, assumption (7) and that $\beta_{N}=o(\kappa_{N})$ in the last line. Step 2. To compute the contribution of binary branch points, we also apply the many-to-few formula, but with a different branch length distribution. We assume that $H^{(N)}_{j}$ are independent and identically distributed with
@@
To distinguish the distribution of this $k$-spine tree from that derived with uniform branch lengths, let us denote it by ¯$\bar{\mathbb{Q}}^{k,Nt}$. We obtain by Proposition 5.3 that
@@
By Lemma A.4 the distribution $H^{(N)}_{i}/N$ under ¯$\bar{\mathbb{Q}}^{k,Nt}$ converges to the law with cumulative distribution function $F$ defined in (16). We want to derive the almost sure limit of $\Delta_{k}$ under ¯$\bar{\mathbb{Q}}^{k,Nt}$ and use the bounded convergence theorem to conclude the proof. Since $t$ is a continuity point of $(X_{s},s\geq 0)$, by Lemma A.4,
@@
¯$\bar{\mathbb{Q}}^{k,Nt}\Big{[}\mathbbm{1}_{|B|<k-1}\Big{]}=O\Big{(}\frac{1}{\kappa_{N}^{k-1}}\bar{\mathbb{Q}}^{k,Nt}\Big{[}\Delta_{k}\mathbbm{1}_{|B|<k-1}\Big{]}\Big{)}=O\Big{(}\frac{1}{\kappa_{N}^{k-1}}\mathbb{E}\Big{[}\sum_{\begin{subarray}{c}u_{1},\dots,u_{k}\in T_{N}\\ \text{$(u_{i})$ distinct}\end{subarray}}\mathbbm{1}_{|B|<k-1}\Big{]}\Big{)}\to 0.$
@@ 7.2 Truncating the offspring distribution
By assumption (6) the second term in (56) converges to $(X_{s},s\geq 0)$. Therefore, we only need to prove that the first term converges to 00 uniformly. Note that $\tilde{f}_{j}^{\prime}(1)\leq f_{j}^{\prime}(1)$, hence
$\left|\sum_{j=1}^{sN}\log\frac{\tilde{f}_{j}^{\prime}(1)}{f_{j}^{\prime}(1)}\right|$ $=\sum_{j=1}^{sN}\log\frac{f_{j}^{\prime}(1)}{\tilde{f}_{j}^{\prime}(1)}\leq\sum_{j=1}^{sN}\frac{f_{j}^{\prime}(1)}{\tilde{f}_{j}^{\prime}(1)}-1$ (57)
@@
Point (i) is thus proved if we can lower bound $\tilde{f}^{\prime}_{j}(1)$ uniformly in $N$ and $j\leq sN$. Write
$\tilde{f}^{\prime}_{j}(1)=f^{\prime}_{j}(1)-(f^{\prime}_{j}(1)-\tilde{f}^{\prime}_{j}(1))\geq f^{\prime}_{j}(1)-\Big{(}\sum_{i=1}^{sN}(f^{\prime}_{i}(1)-\tilde{f}^{\prime}_{i}(1))\Big{)}.$
The second term is going to 00 by (59), whereas $f^{\prime}_{j}(1)$ is uniformly bounded away from 00 by Lemma A.3, hence we obtain the required bound. For point (ii), notice
@@
Consider a sequence of BPVEs verifying (6), and a sequence $(\beta_{N},N\geq 1)$ fulfilling (51). Let $\tilde{Z}_{tN}$ be the size of the BPVE at time $tN$ whose offspring law is truncated by $\beta_{N}$. Then
@@ 7.3 Proof of the Gromov-weak convergence
We now prove that the rescaled genealogy converges in the Gromov-weak topology. Recall that $[T_{tN},d_{tN},\lambda_{tN}]$ is the random metric measure space representing the genealogy of the population at generation $tN$.
@@
with $[(0,Z_{e}),d_{\nu_{e}},\operatorname{Leb}]$ the environmental CPP as defined in (30).
@@
By Lemma A.2 there exists a sequence $\beta_{N}=o(\kappa_{N})$ fulfilling (50). Consider the corresponding truncated BPVE $(\tilde{Z}_{n},n\geq 1)$, and the corresponding metric measure space $[\tilde{T}_{tN},\tilde{d}_{tN},\tilde{\lambda}_{tN}]$ giving the genealogy at time $tN$. By Lemma 7.3 we have that for any $\varepsilon>0$,
@@
Applying Corollary 3.4 to $[T_{tN},\tfrac{d_{tN}}{N},\tfrac{\lambda_{tN}}{\kappa_{N}}]$ and $[\tilde{T}_{tN},\tfrac{\tilde{d}_{tN}}{N},\tfrac{\tilde{\lambda}_{tN}}{\kappa_{N}}]$ under the measure $\mathbb{P}(\>\cdot\mid Z_{tN}>0)$ shows that the result is proved if we can prove that $[\tilde{T}_{tN},\tfrac{\tilde{d}_{tN}}{N},\tfrac{\tilde{\lambda}_{tN}}{\kappa_{N}}]$ converges to the environmental CPP. Up to replacing $T_{tN}$ by $\tilde{T}_{tN}$, we can therefore assume that the environment fulfils (50). By Proposition 3.1 we need to show the following convergence for any polynomial $\Phi$,
@@
where $(H_{i,j})_{i,j}$ are given by (18) and $\sigma\colon[k]\to[k]$ is an independent random permutation of $[k]$. Therefore, we show that the left hand side in (61) converges to the above limit. Using the many-to-few formula, we have shown in Proposition 7.1 that
@@ 7.4 Completing the proof of Theorem 2
$\forall k\geq 0,\quad\mathbb{P}(K^{\prime}=k)=\frac{1}{1+c}\Big{(}\frac{c}{1+c}\Big{)}^{k}$
with $c$ the constant defined in (68). By the branching property, the number of descendants of each of these $Z_{sN,tN}$ individuals is distributed as $Z_{tN}$ conditional on $\{Z_{sN}=1,Z_{tN}>0\}$. Thus, another application of Theorem 3 shows that
@@
We apply Proposition 3.6. We only need to check that (27) holds for the sequence of random metric measure spaces $([T_{tN},\tfrac{d_{tN}}{N},\tfrac{\lambda_{tN}}{\kappa_{N}}])_{N}$. In the context of the BPVE, the balls of radius $t-s$, with $s<t$, correspond to the ancestral lineages at time $s$. That is, each ball corresponds to one of the $Z_{sN,tN}$ individuals at time $sN$ having descendants at $tN$. Moreover, the mass of a ball is the number of descendants of the corresponding ancestor, rescaled by $\kappa_{N}$. In the notation of Corollary 7.4, this corresponds to the random variables $(\tfrac{|T_{tN}|(i)}{\kappa_{N}},i\leq Z_{sN,tN})$. By Corollary 7.4, the number of such balls converges to a geometric random variable, and the mass of the balls converge to i.i.d. exponential random variables. This readily shows that condition (27) is fulfilled, proving the result. ∎
@@
Both maps
$[X,d,\nu]\mapsto\nu(X),\qquad[X,d,\nu]\mapsto[X,d,\tfrac{\nu}{\nu(X)}]$
are continuous w.r.t. the Gromov-weak topology. Since the BPVE converges in the GHP topology by Theorem 2 this readily shows (i), and since for any functional $\phi$, the expectation $\mathbb{E}\big{[}\phi\big{(}\big{(}\tfrac{d_{tN}(U_{i},U_{j})}{N}\big{)}_{i,j}\big{)}\big{]}$ is nothing but the moment of order $k$ of $[(0,Z_{e}),d_{\nu_{e}},Z_{e}^{-1}\operatorname{Leb}]$ associated to $\phi$, Proposition 4.4 proves (iii). The finite-dimensional part of point (ii) follows from Proposition 3.5. For the Skorohod part, note that if $\sigma^{2}$ is continuous, the reduced process associated to the corresponding environmental CPP only makes jump of size $1$ almost surely, and the result follows again from Proposition 3.5. ∎
@@ Appendix A Appendix
We prove the result by induction on $k$. For $k=1$ there is no branch point and the claim holds. By the CPP construction, a uniform ultrametric tree $S_{k+1}$ with $k+1$ leaves is obtained by grafting a branch $k+1$ with uniform length to the right of a uniform ultrametric tree $S_{k}$ with $k$ leaves. There are two options for $S_{k+1}$ to have $b$ branch points. Either $S_{k}$ had $b-1$ branch points and a new branch point is created. Or $S_{k}$ has $b$ branch points, and no new branch point is created. This requires that $H_{k}^{(N)}$ takes the same value as one of the variables $(H_{1}^{(N)},\dots,H^{(N)}_{k-1})$, which occurs with probability at most $(k-1)/N$. Therefore,
@@
We simply adapt the argument of [26, Lemma 22] to our case. By assumption (8), for any fixed $\varepsilon$
@@
It is easily seen that (i) and (ii) are equivalent. We thus only prove (ii). By Lemma A.3, the expectations $f^{\prime}_{k}(1)$ are bounded uniformly in $N$ and $k\leq Nt$, so that (77) yields that for any $\varepsilon>0$ there is $K$ such that

### v4

@@ 7 Convergence of the BPVE
This last section is dedicated to the proof of Theorem 2. Recall the encoding of generation $tN$ of the BPVE as a random metric measure space $T_{tN},d_{tN},\lambda_{tN}]$ from Section 7.
As alluded to in the introduction, we follow the general proof strategy proposed in an earlier work by two of the authors [15]. We start by proving that the genealogy of the BPVE converges in the Gromov-weak topology. According to Proposition 3.1, proving convergence in distribution for the Gromov-weak topology amounts to computing the limit of the moments of the random metric measure space $T_{tN},d_{tN},\lambda_{tN}]$. Using the many-to-few formula, see Proposition 5.3, these moments can be expressed in terms of the distribution of a finite tree, the $k$-spine tree [27, 25, 15]. This $k$-spine tree is studied in Section 7.1 where we derive the limit of the moments of the BPVE. A small caveat to this approach is that, for the moments to be well-defined, we need to make a truncation of the reproduction laws, which is carried out in Section 7.2. We relate the truncated and the untruncated versions via Corollary 3.4. Once the convergence is established in the Gromov-weak topology, we reinforce it to a convergence in the GHP topology by using Proposition 3.6.
@@ 7.1 Moments of the BPVE
Consider a BPVE satisfying (6) and (7), and a fixed $t>0$ such that $t$ is a continuity point of $\sigma^{2}$ and $X$, and $\sigma^{2}$ is strictly increasing at $t$. If the moment growth assumption (47) is fulfilled, then, for any bounded continuous $\phi$
@@
where $(H_{i,j})_{i,j}$ are as in (31) and $\sigma$ is an independent and uniform permutation of $[k]$.
@@
Let $|B_{u_{1},\dots,u_{k}}|$ denote the number of branch points of the subtree spanned by $(u_{1},\dots,u_{k})$. We decompose the expectation with respect to the value of $|B_{u_{1},\dots,u_{k}}|$ as
@@
There are two steps in the proof. First, we prove that, due to our assumption (47) on the moments, the contributions of the terms with b<k1b<k-1 vanish. Second, we compute the limit of the terms with only binary branch points and show that we recover the moments of the environmental CPP.
Step 1. Fix b<k1b<k-1. We apply the many-to-few formula with branch lengths $H_{i}^{(N)}$ uniformly distributed on $\{1,\dots,tN\}$. Let $\mathbb{Q}^{k,Nt}$ be the corresponding $k$-spine tree distribution. By Proposition 5.3, we have
@@
where we have used Lemma A.1 in the second line, assumption (7) and that $\beta_{N}=o(\kappa_{N})$ in the last line.
Step 2. To compute the contribution of binary branch points, we also apply the many-to-few formula, but with a different branch length distribution. We assume that $H^{(N)}_{j}$ are independent and identically distributed with
@@
To distinguish the distribution of this $k$-spine tree from that derived with uniform branch lengths, let us denote it by $\bar{\mathbb{Q}}^{k,Nt}$. We obtain by Proposition 5.3 that
@@
By Lemma A.4 the distribution $H^{(N)}_{i}/N$ under $\bar{\mathbb{Q}}^{k,Nt}$ converges to the law with cumulative distribution function $F$ defined in (16). We want to derive the almost sure limit of $\Delta_{k}$ under $\bar{\mathbb{Q}}^{k,Nt}$ and use the bounded convergence theorem to conclude the proof. Since $t$ is a continuity point of $(X_{s},s\geq 0)$, by Lemma A.4,
@@
$\bar{\mathbb{Q}}^{k,Nt}\Big[\mathbbm{1}_{|B|<k-1}\Big]=O\Big(\frac{1}{\kappa_{N}^{k-1}}\bar{\mathbb{Q}}^{k,Nt}\Big[\Delta_{k}\mathbbm{1}_{|B|<k-1}\Big]\Big)=O\Big(\frac{1}{\kappa_{N}^{k-1}}\mathbb{E}\Big[\sum_{\begin{subarray}{c}u_{1},\dots,u_{k}\in T_{N}\\ \text{$(u_{i})$ distinct}\end{subarray}}\mathbbm{1}_{|B|<k-1}\Big]\Big)\to 0.$
@@ 7.2 Truncating the offspring distribution
By assumption (6), the second term in (53) converges to $(X_{s},s\geq 0)$. Therefore, we only need to prove that the first term converges to 0 uniformly. Note that $\tilde{f}_{j}^{\prime}(1)\leq f_{j}^{\prime}(1)$, hence
$\left|\sum_{j=1}^{sN}\log\frac{\tilde{f}_{j}^{\prime}(1)}{f_{j}^{\prime}(1)}\right|$ $=\sum_{j=1}^{sN}\log\frac{f_{j}^{\prime}(1)}{\tilde{f}_{j}^{\prime}(1)}\leq\sum_{j=1}^{sN}\left(\frac{f_{j}^{\prime}(1)}{\tilde{f}_{j}^{\prime}(1)}-1\right)$ (54)
@@
Point (i) is thus proved if we can lower bound $\tilde{f}^{\prime}_{j}(1)$ uniformly in NN and $j\leq sN$. Write
$\tilde{f}^{\prime}_{j}(1)=f^{\prime}_{j}(1)-(f^{\prime}_{j}(1)-\tilde{f}^{\prime}_{j}(1))\geq f^{\prime}_{j}(1)-\sum_{i=1}^{sN}(f^{\prime}_{i}(1)-\tilde{f}^{\prime}_{i}(1)).$
The second term is going to 0 by (56), whereas $f^{\prime}_{j}(1)$ is uniformly bounded away from 0 by Lemma A.3, hence we obtain the required bound.
For point (ii), notice that
@@
Consider a sequence of BPVEs satisfying (6), and a sequence $(\beta_{N},N\geq 1)$ fulfilling (48). Let $\tilde{Z}_{tN}$ be the size of the BPVE at time tNtN whose offspring law is truncated by $\beta_{N}$. Then
@@ 7.3 Proof of the Gromov-weak convergence
We now prove that the rescaled genealogy converges in the Gromov-weak topology. Recall that $T_{tN},d_{tN},\lambda_{tN}]$ is the random metric measure space representing the genealogy of the population at generation $tN$.
@@
with $(0,Z_{e}),d_{\nu_{e}},\operatorname{Leb}]$ the environmental CPP as defined in (30).
@@
By Lemma A.2 there exists a sequence $\beta_{N}=o(\kappa_{N})$ fulfilling (47). Consider the corresponding truncated BPVE $(\tilde{Z}_{n},n\geq 1)$, and the corresponding metric measure space $\tilde{T}_{tN},\tilde{d}_{tN},\tilde{\lambda}_{tN}]$ giving its genealogy at time $tN$. By Lemma 7.3, for any $\varepsilon>0$,
@@
Applying Corollary 3.4 to $T_{tN},\tfrac{d_{tN}}{N},\tfrac{\lambda_{tN}}{\kappa_{N}}]$ and $\tilde{T}_{tN},\tfrac{\tilde{d}_{tN}}{N},\tfrac{\tilde{\lambda}_{tN}}{\kappa_{N}}]$ under the measure $\mathbb{P}(\>\cdot\mid Z_{tN}>0)$ entails that the result is proved if we can show that $\tilde{T}_{tN},\tfrac{\tilde{d}_{tN}}{N},\tfrac{\tilde{\lambda}_{tN}}{\kappa_{N}}]$ converges to the environmental CPP. Up to replacing $T_{tN}$ by $\tilde{T}_{tN}$, we can therefore assume that the environment fulfils (47).
By Proposition 3.1, we need to show the following convergence, for any polynomial $\Phi$,
@@
where $(H_{i,j})_{i,j}$ are given by (31) and $\sigma\colon[k]\to[k]$ is an independent random permutation of $[k]$. Therefore, we show that the left hand side in (58) converges to the above limit.
Using the many-to-few formula, we have shown in Proposition 7.1 that
@@ 7.4 Completing the proof of Theorem 2
$\forall k\geq 0,\quad\mathbb{P}(K^{\prime}=k)=\frac{1}{1+c}\Big(\frac{c}{1+c}\Big)^{k},$
where the constant $c$ is defined in (63). By the branching property, the number of descendants of each of these $Z_{sN,tN}$ individuals is distributed as $Z_{tN}$ conditional on $\{Z_{sN}=1,Z_{tN}>0\}$. Thus, another application of Theorem 3 shows that, for any continuous bounded $\phi$,
@@
We apply Proposition 3.6. We only need to check that (27) holds for the sequence of random metric measure spaces $([T_{tN},\tfrac{d_{tN}}{N},\tfrac{\lambda_{tN}}{\kappa_{N}}])_{N}$. In the context of the BPVE, the balls of radius tst-s, with $s<t$, correspond to the ancestral lineages at time $s$. That is, each ball corresponds to one of the $Z_{sN,tN}$ individuals at time $sN$ having descendants at $tN$. Moreover, the mass of a ball is the number of descendants of the corresponding ancestor, rescaled by $\kappa_{N}$. In the notation of Corollary 7.4, this corresponds to the random variables $(\tfrac{|T_{tN}|(i)}{\kappa_{N}},i\leq Z_{sN,tN})$. By Corollary 7.4, the number of such balls converges to a geometric random variable, and the mass of the balls converge to i.i.d. exponential random variables. This readily shows that condition (27) is fulfilled, proving the result. ∎
@@
Both maps $X,d,\nu]\mapsto\nu(X)$ and $X,d,\nu]\mapsto[X,d,\tfrac{\nu}{\nu(X)}]$ are continuous w.r.t. the Gromov-weak topology. Since the BPVE converges in the GHP topology by Theorem 2 this readily shows (i), and since for any functional $\phi$, $\mathbb{E}\big[\phi\big(\big(\tfrac{d_{tN}(U_{i},U_{j})}{N}\big)_{i,j}\big)\big]$ is nothing but the moment of order $k$ of $(0,Z_{e}),d_{\nu_{e}},Z_{e}^{-1}\operatorname{Leb}]$ corresponding to $\phi$, Proposition 4.4 proves (iii). The finite-dimensional part of point (ii) follows from Proposition 3.5. For the Skorohod part, note that if $\sigma^{2}$ is continuous, the reduced process of the corresponding environmental CPP only makes jump of size $1$ almost surely, and the result follows again from Proposition 3.5. ∎
@@ Appendix A Appendix
We prove the result by induction on $k$. For $k=1$ there is no branch point and the claim holds. By the CPP construction, a uniform ultrametric tree $S_{k+1}$ with $k+1$ leaves is obtained by grafting a branch $k+1$ with uniform length to the right of a uniform ultrametric tree $S_{k}$ with $k$ leaves. There are two options for $S_{k+1}$ to have $b$ branch points. Either $S_{k}$ had b1b-1 branch points and a new branch point is created. Or $S_{k}$ has $b$ branch points, and no new branch point is created. This requires that $H_{k}^{(N)}$ takes the same value as one of the variables $(H_{1}^{(N)},\dots,H^{(N)}_{k-1})$, which occurs with probability at most (k1)/N(k-1)/N. Therefore,
@@
We simply adapt the argument of [25, Lemma 22] to our case. By assumption (8), for any fixed $\varepsilon$
@@
It is easily seen that (i) and (ii) are equivalent. We thus only prove (ii). By Lemma A.3, the expectations $f^{\prime}_{k}(1)$ are bounded uniformly in NN and $k\leq Nt$, so that (72) yields that for any $\varepsilon>0$ there is $K$ such that